Datum Pty Ltd provided the following fringe benefits for staff members during the FBT year ended 31 March 2019:

  • Alex was provided with (for no cost to Alex) a laptop computer which has a retail selling price of $4,000. This is used by Alex 60% for work related purposes;
  • Linda took home stock with a retail price of $3,100 (she did not pay anything for this stock);

                        Assume Datum Pty Ltd sells identical goods to the public generally.

Required:

Calculate the fringe benefits tax liability for Datum Pty Ltd for the FBT year ended 31 March 2019.
Answer #1

Alex was provided with (for no cost to Alex) a laptop computer which has a retail selling price of $4,000. This is used by Alex 60% for work related purposes and Linda took home stock with a retail price of $3,100 (she did not pay anything for this stock).

Aggregate fringe benefit provided by employer =$4100*60%+3000=$2460+3000=$5460

Assuming the fringe tax rate of 30%

Fringe Benefit taxable=5460/(1-30%)=5460/0.7=$7,800

The tax payable is the fringe benefits taxable amount multiplied by the rate of tax=7800*30%=$2,340

The fringe benefits tax liability for Datum Pty Ltd for the FBT year ended 31 March 2019=$2,340